An aqueous solution is 34.0% by mass ethanol, CH3CH2OH.

The mole fraction of ethanol in the solution is ?


The mole fraction of nickel(II) bromide, NiBr2, in an aqueous solution is 2.00×10-2.

The percent by mass of nickel(II) bromide in the solution is?

The molality of potassium bromide, KBr, in an aqueous solution is 2.44 mol/kg.

What is the mole fraction of potassium bromide in the solution?
Answer #1

1)

Answer

0.1677

Explanation

Consider 100g of the solution

mass of CH3CH2OH = 34.0g

mass of water = 66.0g

Number of moles of CH3CH2OH = 34.0g/46.08g/mol = 0.7378mol

Number of moles of H2O = 66.0g / 18.02g/mol = 3.6626mol

Total number of moles = 0.7378mol + 3.6626mol = 4.4004mol

mole fraction = 0.7378mol/4.4004mol= 0.1677

2)

Answer

19.84%

Explanation

Consider the solution with100g of NiBr2

number of moles of NiBr2 = 100g/218.49g/mol = 0.4577mol

mole fraction of NiBr2 = 0.02

mole fraction of water = 1.00 - 0.02 = 0.98

mole fraction = number of moles / total number of moles

mole fraction of NiBr2 /mole fraction of H2O = number of moles of NiBr2 /number of moles of H2O

0.02/0.98 = 0.4577mol / number of moles of water

Number of moles of water = (0.98/0.02) × 0.4577mol = 22.4273mol

mass of water = 22.4273mol × 18.02g/mol = 404.14g

Total mass = 100g + 404.14g = 504.14g

mass% of NiBr2 = (100g/504.14g)×100 = 19.84%

3)

Answer

0.0421

Explanation

Molality = number of moles of solute per kilogram of solvent

Consider the solution with 1000g of H2O

number of moles of H2O = 1000g/18.02g/mol = 55.494mol

moles of KBr = 2.44mol

Total number of moles = 55.494mol + 2.44mol =

57.934mol

mole fraction of KBr = 2.44mol/ 57.934mol = 0.0421